# Connecting Gravity and Quantum Physics: Primordial Black Holes and the Evolution of the Universe

### v2

@@ IX Formation of Supermassive Black Holes and the Universe’s Primordial Clumping
$R_{H}(Age_{U})=3c\times 8.08\times 10^{60}t_{p}=2.424\times 10^{61}l_{p}$ (73)
Two significant insights can be gained from this unexpected consistency of our results: first, that the suggested method effectively addresses the challenging cosmological problems; and second, that it validates the rapid, non-accelerated expansion of the very early universe ($q=-\frac{\ddot{a}a}{\dot{a}}=2>0$).
@@
The rapid expansion of SMBHs contributes to early cosmic clustering, which aids in the formation of protogalaxies and, subsequently, early galaxies. Observations from the James Webb Space Telescope (JWST) support this scenario, demonstrating that massive and supermassive black holes act as gravitational attractors, facilitating the formation of large cosmic structures in the early universe.
@@ X Baryogenesis driven by PBH Evolution
In the high-energy environment of the early universe, neutrons and protons could freely convert into one another due to the abundant energy available [93, 94]. However, as the universe expanded and cooled, the energy levels dropped, particularly when positron-electron ($e^{+}-e^{-}$) annihilation became significant and neutron decay to protons became more prevalent, as lower available energy could no longer support reverse conversions [79]. In this context, PBHs start to play a role in contributing to ordinary matter through decay processes [5, 1, 55]. The decay interactions of PBHs can be represented as follows:
@@
where $E(n^{0})=0.9396GeV$ is the rest energy of the neutron, $E(c_{q})=1.25GeV$ energy of the free “charm” quark, next in this line after neutron. This constraint implies that PBHs predominantly generate neutrons up to a specific quantum energy level $n_{s}$. The highest limit, $E(c_{q})$, ensures that the PBHs won’t have enough energy to produce charm quarks, restricting particle production to neutrons and lighter particles. By limiting $n_{s}$ to the specified range, the model ensures that the PBH decays predominantly yield baryons, such as protons, rather than other particles that would not enhance baryon asymmetry. The concept assumes that gravitational effects or PBH decay can induce CP violation, essential for generating a matter-antimatter imbalance. PBHs create conditions in which gravitational interactions may inherently result in such violations, eliminating the necessity for alternative CP-violating mechanisms. Here we need the “charm” quark rest energy in order to determine the quantum number $n_{s}(c_{q})$ at which neutron production stops. Using a simple equation
@@
According to Eq. (26) the formation of baryonic matter takes place over roughly $87$ thousand years, while the complete cycle of PBH evaporation and subsequent extinction lasts approximately $190$ thousand years. During the intermediate period, from $87$ to $190$ thousand years, PBHs may continue decaying and emitting radiation, even though baryonic matter production has largely ceased. This radiation increases the universe’s overall radiation density but does not significantly affect the baryon-to-dark matter ratio, which was established during the initial baryonic production phase [95].
@@ XI The Evolution of Dark Matter and Its Impact on Cosmic Structure
It is important to note that, in the initial epochs of the universe, galaxies were closely clustered, forming a substantial “universal supercluster” filled with matter. As the universe expanded rapidly, fractures and separations emerged, leading to the formation of the massive cosmic structures we observe today—immense voids, superclusters, great filaments, and walls. The important features arise as natural consequences of this rapid expansion, logically derived from the previously mentioned results.
@@ XII Conclusion
Our article examines PBHs as crucial factors in the universe’s evolution, specifically the early universe–relating to the Planck scale physics, particularly in relation to quantized gravity and cosmic events. The suggested model links the quantization of black hole characteristics with fundamental physics and cosmology, providing insights into high-energy astrophysics, and cosmic structure formation. This concept achieves the unification of quantum physics and gravity by the groundbreaking discovery of condensed light, namely, the self-gravitating photon Bose-Einstein condensate. This technique is fundamentally based on a geometric understanding that connects the Compton wavelength of the photon condensate to the gravitational radius of PBHs. This connection elucidates a range of quantized attributes—encompassing geometrical, gravitational, thermodynamic, informational, and radiative properties—illuminating the fundamental principles that dictate the formation, development, and ultimate extinction of successive generations of PBHs.
This advancement unveiled new aspects of Planck-scale and PBH physics, facilitating the creation of efficient and robust mathematical instruments. These tools have facilitated accurate computations and resolutions to various fundamental issues in physics and cosmology, encompassing the beginning of time, Big Bang cosmology, the genesis and development of dark matter and dark energy, as well as the mystery of baryogenesis and baryonic asymmetry. Our model also examines the hyper-exponential growth of SMBHs, the birth of primordial galaxies, the evolution of large-scale structures, and the Universe’s two-phase accelerated expansion. Furthermore, it offers very plausible solutions for persistent problems such as the horizon problem, the cosmological constant problem, the information loss conundrum, and the core-cusp dilemma.

### v3

@@ IX Formation of Supermassive Black Holes and the Universe’s Primordial Clumping
$R_{H}(Age_{U})=3c\times 8.08\times 10^{60}t_{p}=2.424\times 10^{61}l_{p}$
Two significant insights can be gained from this unexpected consistency of our results: first, that the suggested method effectively addresses the challenging cosmological problems; and second, that it validates the rapid, non-accelerated expansion of the very early universe ($q=-\frac{\ddot{a}a}{\dot{a}}>0$).
@@
The rapid expansion of SMBHs contributes to early cosmic clustering, which aids in the formation of protogalaxies and, subsequently, early galaxies. Observations from the James Webb Space Telescope (JWST) support this scenario, demonstrating that massive and supermassive black holes act as gravitational attractors, facilitating the formation of galaxies and large cosmic structures in the early universe.
@@ X Baryogenesis driven by PBH Evolution
In the high-energy environment of the early universe, neutrons and protons could freely convert into one another due to the abundant energy available [94, 95]. However, as the universe expanded and cooled, the energy levels dropped, particularly when positron-electron ($e^{+}-e^{-}$) annihilation became significant and neutron decay to protons became more prevalent, as lower available energy could no longer support reverse conversions [80]. In this context, PBHs start to play a decisive role in the production of ordinary matter through decay processes.
The productive decay of PBHs can be represented as follows:
@@
where $E(n^{0})=0.9396GeV$ is the rest energy of the neutron, $E(c_{q})=1.25GeV$ energy of the free “charm” quark, next in this line after neutron. This constraint implies that PBHs predominantly generate neutrons up to a specific quantum energy level $n_{s}$. The highest limit, $E(c_{q})$, ensures that the PBHs won’t have enough energy to produce charm quarks, restricting particle production to neutrons and lighter particles. By limiting $n_{s}$ to the specified range, the model ensures that the PBH decays predominantly yield baryons, such as protons, rather than other particles that would not enhance baryon asymmetry. The concept assumes that gravitational effects or PBH decay can induce CP violation, essential for generating a matter-antimatter imbalance. PBHs create conditions in which gravitational interactions may inherently result in such violations, eliminating the necessity for alternative CP-violating mechanisms. Here we need the “charm” quark rest energy in order to determine the quantum number $n_{s}(c_{q})$ at which neutron production stops. Using a simple equation [Borsevici3]
@@
According to Eq. (37) the formation of baryonic matter takes place over roughly $87$ thousand years, while the complete cycle of PBH evaporation and subsequent extinction lasts approximately $190$ thousand years. During the intermediate period, from $87$ to $190$ thousand years, PBHs may continue decaying and emitting radiation, even though baryonic matter production has largely ceased. This radiation increases the universe’s overall radiation density but does not affect the baryon-to-dark matter ratio, which was established during the initial baryonic production phase [96], in strict conformity with the baryon number conservation law.
@@
These results are important for preserving the well-deserved reputation of the standard $\Lambda$CDM cosmological model by freeing it from the arbitrarily introduced so-called ”hierarchical” models, that contradict JWST data.
@@ XI The Evolution of Dark Matter and Its Impact on Cosmic Structure
It is important to note that, in the initial epochs of the universe, galaxies were closely clustered, forming a substantial ”universal supercluster” filled with dark and baryonic matter. As the universe expanded rapidly, fractures and separations emerged, leading to the formation of the massive cosmic structures we observe today—immense voids, superclusters, great filaments, and walls. The important features arise as natural consequences of this rapid expansion, logically derived from the previously mentioned results.
@@ XII Conclusion
Our article examines PBHs as crucial factors in the universe’s evolution, specifically the early universe–relating to the Planck scale physics, particularly in relation to quantized gravity and cosmic events. The suggested model links the quantization of black hole characteristics with fundamental physics and cosmology, providing insights into high-energy astrophysics, and cosmic structure formation. This concept achieves the unification of quantum physics and gravity by the groundbreaking discovery of condensed light, namely, the self-gravitating photon Bose-Einstein condensate. This technique is fundamentally based on a geometric understanding that connects the Compton wavelength of the photon condensate to the gravitational radius of PBHs. This connection elucidates a range of quantized attributes-encompassing geometrical, gravitational, thermodynamic, informational, and radiative properties-illuminating the fundamental principles that dictate the formation, development, and ultimate extinction of successive generations of PBHs.
This advancement unveiled new aspects of Planck-scale and PBHs physics, facilitating the creation of efficient and robust mathematical instruments. These algebraic tools, coined by A. Einstein in his last lines, have facilitated accurate computations and resolutions to various fundamental issues in physics and cosmology, encompassing the Beginning of Time, Big Bang cosmology, the genesis and development of dark matter and dark energy, as well as the mystery of baryogenesis and baryonic asymmetry. Our model also examines the hyper-exponential growth of SMBHs, the birth of primordial galaxies, the evolution of large-scale structures, and the Universe’s two-phase accelerated expansion.
Furthermore, we offer a very plausible solutions for multiple persistent issues, problems, paradoxes, and ”mysteries” such as cosmological constant problems, informational loss paradox, cosmological principle problems related to the large-scale structure ”excesses”, physical nature and origin of dark matter and dark energy, baryogenesis and baryonic assymetry, the very early growth of SMBHs as the gravitational centers of the very early proto-galaxies, the horizon problem, the core-cusp dilemma, the very early clumping of the Universe and it’s accelerated evolution from the Beginning of Time.